# Flavor solitons in dense neutrino gases

### v1

@@ D.1 From Three Beams to Pendulum and Back: Systematic Lax Vector Approach
The length of the vector ${\bf M}_{1}$ is conserved, so we may write ${\bf M}_{1}=|{\bf M}_{1}|{\bf r}$, with ${\bf r}$ a unit vector, which will be our pendulum direction. We also introduce the definition ${\bf J}={\bf M}_{2}-s{\bf M}_{1}$. Finally, we write ${\bf M}_{0}=|{\bf M}_{0}|\hat{{\bf z}}$, fixed along the $z$-axis. Therefore, the relevant EOMs become
@@ D.2 From Continuous Spectrum to Three Beams: Systematic Matching Conditions
Incidentally, these conditions can be given a simple physical interpretation using an electrostatic analogy (see also Ref. Pehlivan:2011hp ). If we interpret $vD^{z}_{v}$ as a continuous linear charge density, the quantity appearing in the second and third Eqs. (D.2) is the $x$ and $y$ component of the electrostatic field generated at the point $(\mathrm{Re}(\bar{u}),\mathrm{Im}(\bar{u}))$. The charge distribution $vD^{z}_{v}$ vanishes at two points, namely at $v=0$ and $v=v_{c}$, the crossing point of $D^{z}_{v}$. Therefore, the three-beam mapping corresponds to replacing the continuous distribution, split into three intervals with alternating charges, by three effective charges $u_{i}L^{z}_{i}$ placed at the positions $u_{i}$ giving the same total charge, and with an electrostatic field vanishing at the same off-axis point. These effective charges can of course be chosen in the interval $-1<u_{i}<1$, since they represent some kind of average properties of the continuous distribution in the same interval.
@@
If we regard now $\rho_{v}=D^{z}_{v}/|v+\bar{u}|^{2}$ as a distribution over the interval $[-1,1]$, this form shows that the three beams are effectively replacing the continuous distribution with a discrete distribution whose moments of order up to $3$ are identical.

### v2

@@ D.1 From Three Beams to Pendulum and Back: Systematic Lax Vector Approach
The length of the vector ${\bf M}_{1}$ is conserved, so we may write ${\bf M}_{1}=|{\bf M}_{1}|{\bf r}$, with $\bf r$} a unit vector, which will be our pendulum direction. We also introduce the definition ${\bf J}={\bf M}_{2}-s{\bf M}_{1}$. Finally, we write $^{\bf M}_{0}=|{\bf M}_{0}|\hat{{\bf z}}$, fixed along the $z$-axis. Therefore, the relevant EOMs become
@@ D.2 From Continuous Spectrum to Three Beams: Systematic Matching Conditions
Incidentally, these conditions can be given a simple physical interpretation using an electrostatic analogy (see also Ref. Pehlivan:2011hp ). If we interpret $vD^{z}_{v}$ as a continuous linear charge density, the quantity appearing in the second and third Eqs. (110) is the $x$ and $y$ component of the electrostatic field generated at the point $(\mathrm{Re}(\bar{u}),\mathrm{Im}(\bar{u}))$. The charge distribution $vD^{z}_{v}$ vanishes at two points, namely at $v=0$ and $v=v_{c}$, the crossing point of $D^{z}_{v}$. Therefore, the three-beam mapping corresponds to replacing the continuous distribution, split into three intervals with alternating charges, by three effective charges $u_{i}L^{z}_{i}$ placed at the positions $u_{i}$ giving the same total charge, and with an electrostatic field vanishing at the same off-axis point. These effective charges can of course be chosen in the interval $1<u_{i}<1$, since they represent some kind of average properties of the continuous distribution in the same interval.
@@
If we regard now $\rho_{v}=D^{z}_{v}/|v+\bar{u}|^{2}$ as a distribution over the interval [1,1][-1,1], this form shows that the three beams are effectively replacing the continuous distribution with a discrete distribution whose moments of order up to $3$ are identical.